Solve y? + 169 = 0, y(5) = 3, y ( 1 ) = 8 g(t) = Preview
Answer #1

Solution. Given yll+164=0 YI)=3, yC I J = 8. Auxilliary eavation. m2 +16=0 » (m+48)(m-46)=0 - me =-40 m2=4 è be the mot of A